OpenStudy (anonymous):

why is there no infinite geometric series with a first term 10 and a sum 4?

OpenStudy (inkyvoyd):

I"m sure there is.

OpenStudy (anonymous):

there isn't :( the problem asks you to explain why

OpenStudy (inkyvoyd):

Then, I don't know what I'm talking about.

OpenStudy (inkyvoyd):

@Callisto , @saifoo.khan , help pwease?

OpenStudy (saifoo.khan):

No idea.. sorry.. im stuck with "why" as well

OpenStudy (inkyvoyd):

I script kiddied wikipedia, and I know why now.

OpenStudy (inkyvoyd):

S=a/(1-r)

OpenStudy (inkyvoyd):

r must be less than 1, or it doesn't converge.

OpenStudy (callisto):

Assume it exists.. a ---- = Sum to infinity 1-r 10 ---- = 4 1-r 10 = 4-4r 6= -4r r= -3/2 However, for sum of infinity to exist, -1<r<1 . In this case, r=-3/2 < -1 So, it doesn't exist. I don't know if this is correct though :(

OpenStudy (inkyvoyd):

S=4, and r=10

OpenStudy (inkyvoyd):

4-4r=10 6=-4r r=-3/2

OpenStudy (inkyvoyd):

That's what wikipedia says, at least.

OpenStudy (callisto):

I just searched it in my dearest 'last minute' book :(

OpenStudy (anonymous):

no that actually makes a ton of sense! thanks a lot guys!